OpenStudy (anonymous):

Show that the average speed of Leslie is 10 km/h when she runs to the store 5km away in 30 minutes. How fast is that in m/s?

OpenStudy (vijeya3):

The average speed is given by Total distance/total time To convert minutes into hours, 30 minutes=30/60 hours 30 minutes=0.5 hours speed=5km/0.5 hours =10km/h To convert km/h into m/s we multiply the speed in km/h by 5/18, so.. 10km/h*5/18 =2.77 m/s(approx)